OpenStudy (anonymous):

Solve 5x2 = -30x - 65. x = -3 ± 2i x = -3 ± 4i x = -6 ± 2i x = -6 ± 4i help please

OpenStudy (campbell_st):

rewrite the equation as \[ 5x^2+30x + 65 = 0\] divide every term by 5 \[x^2 + 6x + 13 = 0\] then use the general quadratic formula \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] from your equation you need to use a 1, b = 6 and c = 13 substitute and then simplify to find the answers